OpenStudy (anonymous):

Help!! A Line having a slope of 3/4 passes through the point (-8,4). Write the equation of this line in slope-intercept form.

OpenStudy (anonymous):

See here you have slope and one point also so use this formula: \[y = y_1 = m(x - x_1)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[y - y_1 =m(x - x_1)\]

OpenStudy (anonymous):

wait i don't get it :(

OpenStudy (anonymous):

Here: \(y_1 = 4\) and \(x_1 = -8\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This is one point slope form: \[y - (-8) = \frac{3}{4}(x - 4)\]

OpenStudy (anonymous):

As slope (m) = 3/4 so I have just plugged in the values in place of x1, y1 and m..

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

So just solve that equation..

OpenStudy (anonymous):

omg i'm so lost :(( lol

OpenStudy (anonymous):

\[y + 8 = \frac{3}{4}(x - 4) \implies 4(y + 8) = 3(x - 4) \implies 4y + 32 = 3x - 12\] \[4y = 3x - 12 - 32 \implies 4y = 3x - 44\] Now divide by 4 and you will have you answer..

OpenStudy (anonymous):

how did you get 4?

OpenStudy (anonymous):

Where??

OpenStudy (anonymous):

y+8=3/4(x-4) to 4(y+8) ? like how did you get it.

OpenStudy (anonymous):

Multiply by 4 both the sides...

OpenStudy (anonymous):

thank you :)))

OpenStudy (anonymous):

You are welcome dear..

OpenStudy (anonymous):

:)